OpenStudy (anonymous):

express the repeating decimal 0.23232323... as a geometric series, and write its sum as the ratio of two integers.

OpenStudy (anonymous):

\[.232323...=23\times (\frac{1}{100}+\frac{1}{100^2}+\frac{1}{100^3}+...\] is one way to write is as a geometric series

OpenStudy (anonymous):

not to hard to solve for the fraction, even without using a geometric series \[x=.232323...\\ 100x=23.232323...\\99x=23\\x=\frac{23}{99}\]

OpenStudy (anonymous):

or you can use \[23\times \frac{\frac{1}{100}}{1-\frac{1}{100}}\] \[=23\times \frac{\frac{1}{100}}{\frac{99}{100}}\] \[=23\times \frac{1}{99}\] \[=\frac{23}{99}\] either way

OpenStudy (anonymous):

or you can remember that if it is a repeating decimal where two digits repeat, it is that number over 99

OpenStudy (anonymous):

always?

OpenStudy (anonymous):

if it is something like \[.121212...\] then yes, you get \[\frac{12}{99}=\frac{4}{33}\] you can check it with a calculator

OpenStudy (anonymous):

if 3 digits repeat, divide by 999 like \[.123123123...=\frac{123}{999}\]

OpenStudy (anonymous):

if one digit repeats, divide by 9 for example \[.3333...=\frac{3}{9}=\frac{1}{3}\]

OpenStudy (anonymous):

why 99??

OpenStudy (anonymous):

the short answer is because we write out decimals in base ten the long answer is that is what you get when you sum the geometric series see the calculation above and see that you get \(\frac{1}{99}\)

OpenStudy (anonymous):

if that is not a satisfactory answer, we can do it slowly and see why it is true

OpenStudy (anonymous):

okay so, that is how you write down the sum of the ratios? each power increases by 1?

OpenStudy (anonymous):

\[.23=\frac{23}{100}\\.0023=\frac{23}{10000}=\frac{23}{100^2}\\ .000023=\frac{23}{100^3}\] etc

OpenStudy (anonymous):

the font is so weird, your 3 looks like a two haha. okay i see now!

OpenStudy (anonymous):

factor out the \(23\) and the remaining sum is \[\frac{1}{100}+\frac{1}{100^2}+\frac{1}{100^3}+...\] if you are familiar with sigma notation, it is the same as \[\sum_{k=1}^{\infty}(\frac{1}{100})^k\]

OpenStudy (anonymous):

yeah it is large on my screen but i have heard that before

OpenStudy (anonymous):

oooh okay i see.

OpenStudy (anonymous):

oh haha yah it looks normal now that i made my screen bigger lol.

OpenStudy (anonymous):

would i write the 23 outside of the (1/100)?